# Edward Witten

### v1

@@ 2 Black Hole Entropy And The Generalized Second Law
In $3+1$ dimensions, [^5] We generally phrase arguments in this article for the standard $3+1$-dimensional case, as the generalization to $D$ dimensions would not add much. The discussion of the Ryu-Takayanagi formula is the main exception, for reasons explained in footnote 48. the relation between energy $E$, temperature $T$, and entropy $S$ in black body radiation is $E=\frac{3}{4}TS$. If a black hole of mass $M$ absorbs energy $E\ll M$ from black body radiation, its entropy $A/4G=4\pi GM^{2}$ increases by $8\pi GME$, while the entropy of the radiation decreases by $\Delta S_{\rm out}=4E/3T$. The change in the generalized entropy is
@@ 3 Black Hole Evaporation
An important example is a Penrose diagram describing spherically symmetric collapse of a star to a Schwarzschild black hole, as in fig. 1. The left vertical boundary of the picture is the origin of polar coordinates at $r=0$. Shown in red is the worldvolume of the star. (The fact that the star appears to emanate from a point in the far past is an artifact of the conformal mapping used in drawing the diagram.) The star ends its life at the singularity, depicted by the wiggly line at the top of the diagram. Future and past null infinity are represented by the diagonal lines on the right boundary, as labeled. The horizon of the black hole is the diagonal black line inside the picture. Since causal curves travel at an angle no greater than $\pi/4$ from the horizon, an observer outside the horizon can never see beyond the horizon. The worldline of a massive observer who remains forever outside the horizon (and does not accelerate indefinitely) will end at the point ${\rm i}^{+}$, known as future infinity, where the horizon and future null infinity meet. Anything that can be seen from anywhere in the spacetime outside the horizon can be seen from ${\rm i}^{+}$. A massive observer began life in the far past, at the point labeled ${\rm i}^{-}$ at the bottom of the figure.
@@
Let $u$ be any coordinate function on ${\mathcal{S}}$ that vanishes on the horizon and is positive outside, and such that the normal derivative of $u$ is nonzero (and finite) along the horizon. An outgoing signal from the black hole that will eventually be received by a distant observer might originate at any value of $u$. And let $t$ be the time at which the distant observer receives the signal. The relation between $t$ and $u$ is
@@
Setting $u=e^{C}e^{-t/4GM}$, we can turn eqn. (10) into a formula for the two-point function measured by the distant observer at late times:
@@
This is antiperiodic in imaginary time; in fact, it is odd under $t\to t+8\pi GM{\rm i}$. Antiperiodicity with that period corresponds to a thermal correlation function [^9] At temperature $T$, fermion correlation functions are antiperiodic under $t\to t+2\pi{\rm i}/T$, and boson correlation functions are periodic. For a calculation similar to the one in the text with a bosonic field, one could consider a chiral current $J$, of dimension 1, with vacuum expectation value $\langle J(u)J(u^{\prime})\rangle=\frac{{\mathrm{d}}u\,{\mathrm{d}}u^{\prime}}{(u-u^{\prime})^{2}}$. Changing variables from $u$ to $t$, one finds periodicity of $\langle J(t)J(t^{\prime})\rangle$ under $t\to t+8\pi GM{\rm i}$. at a temperature $T_{\rm H}=1/8\pi GM$, known as the Hawking temperature of the black hole. In fact, the right hand side of eqn. (11) is the two-point function of a chiral free fermion at temperature $T_{\rm H}$. In other words it is the thermodynamic limit of $\frac{1}{Z}{\rm Tr}\,e^{-\beta_{\rm H}H}\psi(t)\psi(t^{\prime})$, where $H$ is the Hamiltonian and $Z$ is the partition function of a chiral free fermion, and $\beta_{\rm H}=1/T_{\rm H}$. This statement can be verified by a standard textbook calculation, but such a calculation is not really necessary, since this thermal two-point function is uniquely determined by the following facts: it is antiperiodic under $t\to t+8\pi GM{\rm i}$, and modulo this antiperiodicity, its only singularity is a simple pole at $t=t^{\prime}$ with residue 1 (fig. 3).
@@
The most important application of the Kruskal-Szekeres coordinates is to describe the extension of the Schwarzschild geometry beyond the horizon at $r=2GM$. We will return to this in section 6.7. For now, we will just use these coordinates to justify eqn. (8). The form (16) of the metric shows that a radially outgoing or ingoing null geodesic [^10] A radial null geodesic is defined as one located at a fixed value of the polar angles, so that ${\mathrm{d}}\Omega^{2}=0$ along such a geodesic. Hence ${\mathrm{d}}U\,{\mathrm{d}}V=0$ along such a geodesic and either ${\mathrm{d}}U=0$ or ${\mathrm{d}}V=0$. must satisfy ${\mathrm{d}}U=0$ or ${\mathrm{d}}V=0$, so in other words $U$ or $V$ is constant along such a geodesic. More specifically, from the formulas (14), we see that $U$ is constant on an outgoing radial null geodesic and $V$ is constant on an ingoing one. Eqn. (14) shows that $U$ vanishes at $r=2GM$ and is negative for $r>2GM$, so for a function that vanishes on the horizon and is positive outside, we can take $u=-U$. Eqn. (14) then gives the claimed result $u=Ce^{-t/4GM}$, where $C=\left(\frac{r}{2GM}-1\right)^{1/2}e^{r/4GM}$ is a constant that depends on the position of the observer and not on the time $t$ at which an observation is made. Actually, it is natural to introduce the retarded time $t_{\rm ret}=t-r$ and write the formula for $u$ in the form
@@ 4.1 The Potential Barrier
To get from the horizon to infinity, a wave will have to propagate over the potential barrier. Only a wave whose energy is much greater than $\sqrt{V_{\rm max}(l)}$ will propagate almost freely [^12] Such a wave still experiences a phase shift or time delay, but this is not noticeable to the distant observer, who does not know when the signal was emitted. from the horizon at $r_{*}=-\infty$ across the barrier to $r_{*}=+\infty$. So our previous calculation is good for the $l=0$ mode if the Hawking radiation from the black hole is being observed at frequencies much above the Hawking temperature $T_{\rm H}=\frac{1}{8\pi GM}\sim\sqrt{V_{\rm max}(l=0)}$. For $l>0$, the previous calculation is good at frequencies much above $lT_{\rm max}$. At any given frequency, the analysis of section 3 is completely wrong if $l$ is sufficiently large, because a wave of any given energy is far below the barrier if $l$ is sufficiently large.
@@
Before giving a technical justification of this claim, we will first explain the implications. Suppose that it is true that the black hole has a temperature $T_{\rm H}$. Then we expect that it can be in equilibrium with a thermal gas at that temperature. In equilibrium with such a gas, the black hole is absorbing thermal radiation at temperature $T_{\rm H}$; equilibrium can potentially be maintained because the black hole is also emitting thermal radiation at the same temperature. In the simplest situation, assume that the effective potential vanishes for some partial wave. (As already noted, this only happens for an electrically charged massless fermion interacting with a magnetically charged black hole.) Then the black hole will absorb all of the incident radiation in that partial wave. Equilibrium is maintained because, as analyzed in section 3, in the absence of a potential barrier, the black hole is also freely emitting thermal radiation at temperature $T_{\rm H}$. Now consider a more realistic situation with $V_{\rm eff}\not=0$. Not all incident radiation is absorbed. In a given partial wave at a given energy, the absorption probability is reduced by a factor equal to the transmission probability from right to left (that is, from $r_{*}=+\infty$ to $r_{*}=-\infty$). To maintain equilibrium, the emission probability must be reduced by the same factor. However, as will be explained in a moment, the transmission probability through the barrier from left to right is the same as the transmission probability from right to left. So equilibrium of the black hole with thermal radiation at temperature $T_{\rm H}$ is possible if the emission from the black hole is reduced relative to our previous result by a factor of the transmission probability from left to right through the barrier.
@@ 4.3 Thermodynamic Instability
Given this, what is the sense of the argument that we gave concerning equilibrium between a black hole and a thermal gas? The answer to this question is that, as the derivation of the Hawking effect is based on quantum field theory in a fixed spacetime background, the whole analysis is asymptotically valid in the limit $G\to 0$. As $G\to 0$, the various instabilities that we have mentioned turn off. This is actually visible in eqn. (34). Setting $T$ to be the Hawking temperature $T_{\rm H}=1/8\pi GM$ of a black hole of mass $M$, we find that a thermal gas can occupy a region with radius of order $G^{3/2}M^{2}$. This exceeds the Schwarzschild radius $2GM$ of the black hole by a factor of order $G^{1/2}M$. If this factor is large (it is of order $10^{38}$ for a solar mass black hole), then although the black hole cannot be sensibly embedded in a thermal gas that fills all of an asymptotically flat space, it can be embedded in a thermal gas that fills an enormously large volume.
The instability involving a fluctuation in the mass of the black hole also turns off as $G\to 0$. To see this most clearly, we have to decide what we want to keep fixed as $G\to 0$. There is a factor of $G$ in the relation $r_{S}=2GM$ between the Schwarzschild radius $r_{S}$ and the mass $M$, so we cannot keep them both fixed as $G\to 0$. In studying black holes, it is more natural to keep $r_{S}$ fixed as $G\to 0$. Indeed, in section 3, we worked in a limit $G\to 0$, since we did not consider loop effects due to quantum gravity fluctuations, and we worked in a definite spacetime that had a limit as $G\to 0$, which corresponds to keeping fixed $r_{S}$ rather than $M$ as $G\to 0$. It is natural, then, to express the Hawking temperature in terms of $r_{S}$ rather than $M$: $T_{\rm H}=\frac{1}{4\pi r_{S}}.$ Thus in the limit $G\to 0$, the Hawking temperature is fixed and the black hole mass $M=r_{S}/2G$ diverges. When a statistical fluctuation occurs and the black hole emits a few more particles than it absorbs, its mass drops by an amount of order 1 and its temperature increases by an amount of order $G$. This will indeed cause the black hole to continue to emit more than it absorbs, but only by an amount of order $G$, and the time scale for the resulting instability to significantly change the black hole mass and the associated geometry is of order $1/G^{2}$.
@@ 5.1 Making The Cut
Following are a few important facts about the geometry of Rindler space. First of all, in order to remain for all times in the right Rindler wedge $R_{r}$, an observer must accelerate indefinitely in the future and also in the past, as in eqn. (45) below. (This will lead to the Unruh effect, as we will discuss.) The portions of the spacetime visible to such an observer are bounded by the past and future horizons of Rindler space, the diagonal lines in fig. 7. Similarly, an accelerating observer who remains always in $R_{\ell}$ experiences past and future horizons. The past and future horizons of both observers intersect at a codimension two surface $\Sigma$ known as the bifurcation surface or entangling surface, [^17] In classical general relativity, $\Sigma$ is called the bifurcation surface from which the left and right horizons bifurcate. Quantum mechanically, it is sometimes called the entangling surface of the Rindler wedges $\mathcal{R}_{\ell}$ and $\mathcal{R}_{r}$, which are entangled across $\Sigma$. which also marks the common boundary of ${\mathcal{S}}_{\ell}$ and ${\mathcal{S}}_{r}$. Once we introduce the Penrose diagram of de Sitter space (fig. 10) and the Kruskal-Szekeres extension of Schwarzschild spacetime (fig. 14), it will be clear that all this has an analog for de Sitter space and for a black hole. Indeed, Rindler space is analogous to the near horizon region of a Schwarzschild black hole of mass $M$ in the limit $M\to\infty$, and to the region near the cosmological horizon in de Sitter space. That is the reason for its importance.
@@ 5.4 Another View Of The Thermofield Double
is not holomorphic in the upper half $v$-plane. To get an annihilation operator for the Minkowski vacuum that is equivalent to $U_{\omega},~{}\omega<0$ for observations in the right Rindler wedge, we need to modify $f(v)$ to be non-zero for $v<0$ in such a way that $f(v)$ becomes holomorphic and bounded in the upper half-plane. A function that coincides with $f(v)$ for $v>0$ and is bounded in the upper half $v$-plane is $(v+{\rm i}\epsilon)^{-{\rm i}\omega-\frac{1}{2}}$. Here $\epsilon$ is an infinitesimal positive quantity; a limit $\epsilon\to 0^{+}$ is understood. The boundedness in the upper half plane holds for either sign of $\omega$. Hence, for all $\omega$,
@@ 6.1 Continuing to Euclidean Signature
Fundamentally, the reason that the simple definition $t=-{\rm i}t_{\sf E}$ leads to a real metric in Euclidean signature is the following. The Schwarzschild metric is real, so it has a complex conjugation symmetry $t\leftrightarrow\overline{t}$. This corresponds in terms of $t_{\sf E}$ to $t_{\sf E}\leftrightarrow-\overline{t}_{\sf E}$, which does not imply that the metric is real for real $t_{\sf E}$. But the Schwarzschild spacetime also has a time-reversal symmetry $t\leftrightarrow-t$. The combination of this with complex conjugation is the symmetry $t\leftrightarrow-\overline{t}$, or $t_{\sf E}\leftrightarrow\overline{t}_{\sf E}$, which does indeed imply that the metric is real for real $t_{\sf E}$.

### v4

@@ 2 Black Hole Entropy And The Generalized Second Law
In $3+1$ dimensions, [^5] We generally phrase arguments in this article for the standard $3+1$-dimensional case, as the generalization to $D$ dimensions would not add much. The discussion of the Ryu-Takayanagi formula is the main exception, for reasons explained in footnote 49. the relation between energy $E$, temperature $T$, and entropy $S$ in black body radiation is $E=\frac{3}{4}TS$. If a black hole of mass $M$ absorbs energy $E\ll M$ from black body radiation, its entropy $A/4G=4\pi GM^{2}$ increases by $8\pi GME$, while the entropy of the radiation decreases by $\Delta S_{\rm out}=4E/3T$. The change in the generalized entropy is
@@ 3 Black Hole Evaporation
An important example is a Penrose diagram describing spherically symmetric collapse of a star to a Schwarzschild black hole, as in fig. 1. The left vertical boundary of the picture is the origin of polar coordinates at $r=0$. Shown in green is the worldvolume of the star. (The fact that the star appears to emanate from a point in the far past is an artifact of the conformal mapping used in drawing the diagram.) The star ends its life at the singularity, depicted by the wiggly line at the top of the diagram. Future and past null infinity are represented by the diagonal lines on the right boundary, as labeled. The horizon of the black hole is the diagonal black line inside the picture. Since causal curves travel at an angle no greater than $\pi/4$ from the horizon, an observer outside the horizon can never see beyond the horizon. The worldline of a massive observer who remains forever outside the horizon (and does not accelerate indefinitely) will end at the point ${\rm i}^{+}$, known as future infinity, where the horizon and future null infinity meet. Anything that can be seen from anywhere in the spacetime outside the horizon can be seen from ${\rm i}^{+}$. A massive observer began life in the far past, at the point labeled ${\rm i}^{-}$ at the bottom of the figure.
@@
Let $u$ be any coordinate function on ${\mathcal{S}}$ that vanishes on the horizon and is positive outside, and such that the normal derivative of $u$ is nonzero (and finite) along the horizon. An outgoing signal from the black hole that will eventually be received by a distant observer might propagate through ${\mathcal{S}}$ at any value of $u$. Let $t$ be the time at which the distant observer receives the signal. The relation between $t$ and $u$ is
@@
Setting $u=e^{C/4GM}e^{-t/4GM}$, we can turn eqn. (10) into a formula for the two-point function measured by the distant observer at late times:
@@
This is antiperiodic in imaginary time; in fact, it is odd under $t\to t+8\pi GM{\rm i}$. Antiperiodicity with that period corresponds to a thermal correlation function [^9] At temperature $T$, fermion correlation functions are antiperiodic under $t\to t+{\rm i}/T$, and boson correlation functions are periodic. For a calculation similar to the one in the text with a bosonic field, one could consider a chiral current $J$, of dimension 1, with vacuum expectation value $\langle J(u)J(u^{\prime})\rangle=\frac{{\mathrm{d}}u\,{\mathrm{d}}u^{\prime}}{(u-u^{\prime})^{2}}$. Changing variables from $u$ to $t$, one finds periodicity of $\langle J(t)J(t^{\prime})\rangle$ under $t\to t+8\pi GM{\rm i}$. at a temperature $T_{\rm H}=1/8\pi GM$, known as the Hawking temperature of the black hole. In fact, the right hand side of eqn. (11) is the two-point function of a chiral free fermion at temperature $T_{\rm H}$. In other words it is the thermodynamic limit of $\frac{1}{Z}{\rm Tr}\,e^{-\beta_{\rm H}H}\psi(t)\psi(t^{\prime})$, where $H$ is the Hamiltonian and $Z$ is the partition function of a chiral free fermion, and $\beta_{\rm H}=1/T_{\rm H}$. This statement can be verified by a standard textbook calculation, but such a calculation is not really necessary, since this thermal two-point function is uniquely determined by the following facts: it is antiperiodic under $t\to t+8\pi GM{\rm i}$, and modulo this antiperiodicity, its only singularity is a simple pole at $t=t^{\prime}$ with residue 1 (fig. 3).
@@
The most important application of the Kruskal-Szekeres coordinates is to describe the extension of the Schwarzschild geometry beyond the horizon at $r=2GM$. We will return to this in section 6.7. For now, we will just use these coordinates to justify eqn. (8). The form (16) of the metric shows that a radially outgoing or ingoing null geodesic [^10] A radial null geodesic is defined as one located at a fixed value of the polar angles, so that ${\mathrm{d}}\Omega^{2}=0$ along such a geodesic. Hence ${\mathrm{d}}U\,{\mathrm{d}}V=0$ along such a geodesic and either ${\mathrm{d}}U=0$ or ${\mathrm{d}}V=0$. must satisfy ${\mathrm{d}}U=0$ or ${\mathrm{d}}V=0$, so in other words $U$ or $V$ is constant along such a geodesic. More specifically, from the formulas (14), we see that $U$ is constant on an outgoing radial null geodesic and $V$ is constant on an ingoing one. Eqn. (14) shows that $U$ vanishes at $r=2GM$ and is negative for $r>2GM$, so for a function that vanishes on the horizon and is positive outside, we can take $u=-U$. Eqn. (14) then gives the claimed result $u=C^{\prime}e^{-t/4GM}$, where $C^{\prime}=\left(\frac{r}{2GM}-1\right)^{1/2}e^{r/4GM}$ is a constant that depends on the position of the observer and not on the time $t$ at which an observation is made. Actually, it is natural to introduce the retarded time $t_{\rm ret}=t-r$ and write the formula for $u$ in the form
@@ 4.1 The Potential Barrier
To get from the horizon to infinity, a wave will have to propagate over the potential barrier. Only a wave whose energy is much greater than $\sqrt{V_{\rm max}(l)}$ will propagate almost freely [^12] Such a wave still experiences a phase shift or time delay, but this is not noticeable to the distant observer, who does not know when the signal was emitted. from the horizon at $r_{*}=-\infty$ across the barrier to $r_{*}=+\infty$. So our previous calculation is good for the $l=0$ mode if the Hawking radiation from the black hole is being observed at frequencies much above the Hawking temperature $T_{\rm H}=\frac{1}{8\pi GM}\sim\sqrt{V_{\rm max}(l=0)}$. For $l>0$, the previous calculation is good at frequencies much above $lT_{H}$. At any given frequency, the analysis of section 3 is completely wrong if $l$ is sufficiently large, because a wave of any given energy is far below the barrier if $l$ is sufficiently large.
@@
Before giving a technical justification of this claim, we will first explain the implications. Suppose that it is true that the black hole has a temperature $T_{\rm H}$. Then we expect that it can be in equilibrium with a thermal gas at that temperature Equil1 ; Equil2 . In equilibrium with such a gas, the black hole is absorbing thermal radiation at temperature $T_{\rm H}$; equilibrium can potentially be maintained because the black hole is also emitting thermal radiation at the same temperature. In the simplest situation, assume that the effective potential vanishes for some partial wave. (As already noted, this only happens for an electrically charged massless fermion interacting with a magnetically charged black hole.) Then the black hole will absorb all of the incident radiation in that partial wave. Equilibrium is maintained because, as analyzed in section 3, in the absence of a potential barrier, the black hole is also freely emitting thermal radiation at temperature $T_{\rm H}$. Now consider a more realistic situation with $V_{\rm eff}\not=0$. Not all incident radiation is absorbed. In a given partial wave at a given energy, the absorption probability is reduced by a factor equal to the transmission probability from right to left (that is, from $r_{*}=+\infty$ to $r_{*}=-\infty$). To maintain equilibrium, the emission probability must be reduced by the same factor. However, as will be explained in a moment, the transmission probability through the barrier from left to right is the same as the transmission probability from right to left. So equilibrium of the black hole with thermal radiation at temperature $T_{\rm H}$ is possible if the emission from the black hole is reduced relative to our previous result by a factor of the transmission probability from left to right through the barrier.
@@ 4.3 Thermodynamic Instability
Given this, what is the sense of the argument that we gave concerning equilibrium between a black hole and a thermal gas? The answer to this question is that, as the derivation of the Hawking effect is based on quantum field theory in a fixed spacetime background, the whole analysis is asymptotically valid in the limit $G\to 0$. As $G\to 0$, the various instabilities that we have mentioned turn off. To see this most clearly, we have to decide what we want to keep fixed as $G\to 0$. There is a factor of $G$ in the relation $r_{S}=2GM$ between the Schwarzschild radius $r_{S}$ and the mass $M$, so we cannot keep them both fixed as $G\to 0$. In studying black holes, it is more natural to keep $r_{S}$ fixed as $G\to 0$. Indeed, in section 3, we worked in a limit $G\to 0$, since we did not consider loop effects due to quantum gravity fluctuations, and we worked in a definite spacetime that had a limit as $G\to 0$, which corresponds to keeping fixed $r_{S}$ rather than $M$ as $G\to 0$. It is natural, then, to express the Hawking temperature in terms of $r_{S}$ rather than $M$: $T_{\rm H}=\frac{1}{4\pi r_{S}}.$ Thus in the limit $G\to 0$, the Hawking temperature is fixed and the black hole mass $M=r_{S}/2G$ diverges.
With this in mind, let us re-examine the upper bound (34) for the maximum size $R$ of a region of thermal radiation at temperature $T$ that does not collapse to a black hole. Setting $T$ to be the Hawking temperature $T_{\rm H}=1/4\pi r_{S}$, we find that the upper bound $R\leq 16\pi^{2}r_{S}^{2}/\sqrt{G}$ is of order $1/\sqrt{G}$ as $G\to 0$ with fixed $r_{S}$. If the maximum allowed $R$ is much greater than $r_{S}$ (the upper bound on $R/r_{S}$ is of order $10^{38}$ for a solar mass black hole), then although the black hole cannot be sensibly embedded in a thermal gas that fills all of an asymptotically flat space, it can be embedded in a thermal gas that fills an enormously large volume.
The instability involving a fluctuation in the mass of the black hole also turns off as $G\to 0$. When a statistical fluctuation occurs and the black hole emits a few more particles than it absorbs, its mass drops by an amount of order 1 and its temperature increases by an amount of order $G$. This will indeed cause the black hole to continue to emit more than it absorbs, but only by an amount of order $G$, and the time scale for the resulting instability to significantly change the black hole mass and the associated geometry is of order $1/G^{2}$.
@@ 5.1 Making The Cut
Following are a few important facts about the geometry of Rindler space. First of all, in order to remain for all times in the right Rindler wedge $R_{r}$, an observer must accelerate indefinitely in the future and also in the past, as in eqn. (45) below. (This will lead to the Unruh effect, as we will discuss.) The portions of the spacetime visible to such an observer are bounded by the past and future horizons of Rindler space, the diagonal lines in fig. 7. Similarly, an accelerating observer who remains always in $R_{\ell}$ experiences past and future horizons. The past and future horizons of both observers intersect at a codimension two surface $\Sigma$ known as the bifurcation surface or entangling surface, [^17] In classical general relativity, $\Sigma$ is called the bifurcation surface from which the left and right horizons bifurcate. Quantum mechanically, it is sometimes called the entangling surface of the Rindler wedges $\mathcal{R}_{\ell}$ and $\mathcal{R}_{r}$, which are entangled across $\Sigma$. which also marks the common boundary of ${\mathcal{S}}_{\ell}$ and ${\mathcal{S}}_{r}$. Once we introduce the Penrose diagram of de Sitter space (fig. 10) and the Kruskal-Szekeres extension of Schwarzschild spacetime (fig. 14), it will be clear that all this has analogs for de Sitter space and for a black hole. Indeed, Rindler space is analogous to the near horizon region of a Schwarzschild black hole of mass $M$ in the limit $M\to\infty$, and to the region near the cosmological horizon in de Sitter space. That is the reason for its importance.
@@ 5.4 Another View Of The Thermofield Double
is not holomorphic in the upper half $v$-plane. To get an annihilation operator for the Minkowski vacuum that is equivalent to $U_{\omega},~{}\omega<0$ for observations in the right Rindler wedge, we need to modify $f(v)$ to be non-zero for $v<0$ in such a way that $f(v)$ becomes holomorphic and bounded in the upper half-plane. A function that coincides with $f(v)$ for $v>0$ and is holomorphic and bounded in the upper half $v$-plane is $(v+{\rm i}\epsilon)^{-{\rm i}\omega-\frac{1}{2}}$. Here $\epsilon$ is an infinitesimal positive quantity; a limit $\epsilon\to 0^{+}$ is understood. The boundedness in the upper half plane holds for either sign of $\omega$. Hence, for all $\omega$,
@@ 6.1 Continuing to Euclidean Signature
Fundamentally, the reason that the simple definition $t=-{\rm i}t_{\sf E}$ leads to a real metric in Euclidean signature is the following. The Schwarzschild metric is real for real $t$, so if $t$ is regarded as a complex variable, then the Schwarzschild metric is complex conjugated under $t\leftrightarrow\overline{t}$. This corresponds in terms of $t_{\sf E}$ to $t_{\sf E}\leftrightarrow-\overline{t}_{\sf E}$. The fact that this operation complex conjugates the metric does not imply that the metric is real for real $t_{\sf E}$. But the Schwarzschild metric is invariant under the time-reversal symmmetry $t\leftrightarrow-t$, so it is also complex conjugated by the combined action of time-reversal and complex conjugation of $t$. The combined operation is $t\leftrightarrow-\overline{t}$, or $t_{\sf E}\leftrightarrow\overline{t}_{\sf E}$. The fact that this operation complex conjugates the metric does indeed imply that the metric is real for real $t_{\sf E}$.